# Mixed Membership Estimation for Weighted Networks

### v2

@@ Abstract
Community detection in overlapping un-weighted networks in which nodes can belong to multiple communities is one of the most popular topics in modern network science during the last decade. However, community detection in overlapping weighted networks in which edge weights can be any real values remains a challenge. In this article, to model overlapping weighted networks with latent community memberships, we propose a generative model called the degree-corrected mixed membership distribution-free model which can be viewed as generalizing several previous models. First, we address the community membership estimation of the proposed model by an application of a spectral algorithm and establish a theoretical guarantee of consistency. We then propose overlapping weighted modularity to measure the quality of overlapping community detection for weighted networks with positive and negative edge weights. To determine the number of communities for weighted networks, we incorporate the algorithm into the overlapping weighted modularity. We demonstrate the advantages of degree-corrected mixed membership distribution-free model and overlapping weighted modularity with applications to simulated data and eleven real-world networks.
@@
[…]
[label]organization=School of Mathematics, China University of Mining and Technology,city=Xuzhou, postcode=221116, state=Jiangsu, country=China
@@ 1 Introduction
Community detection is a popular data mining problem in network science and it is to partition a network into several components, where each component is called a community (also referred to as a module or a cluster, or a block). Generally, nodes within the same community have more interactions than nodes across communities. Community detection has many real-world applications in various domains, like sociology, biology, marketing, and computer science girvan2002community ; newman2003the ; barabasi2004network ; guimera2005functional ; palla2007quantifying ; andersen2012overlapping ; garcia2018applications . For example, community detection in biological networks helps researchers to analyze the interaction of brain regions and have a better understanding of the influence of the interaction on brain functions garcia2018applications ; huang2019ocomm ; tooley2022functional . Community detection in social networks can help to identify crime organizations magalingam2015using ; legramanti2022extended . Community detection in online social networks and food social networks can facilitate recommender systems sahebi2011community ; gasparetti2021community ; rostami2023novel . Community detection in coauthorship and citation networks can be used to identify meaningful communities for scientists Newman2004 ; ji2022co .
@@
For its wide applications, in past decades, substantial methods have been developed to learn community structure for non-overlapping un-weighted networks in which each node belongs to a single community and the edges or links between nodes are either present or not fortunato2010community ; fortunato2016community ; papadopoulos2012community , where un-weighted networks are also referred to as binary networks lu2014algorithms . For community detection in non-overlapping un-weighted networks, perhaps the most popular and well-studied generative model is the Stochastic Blockmodel (SBM) SBM . The SBM assumes that the probability of generating an edge between two nodes depends on their belonging communities. Under SBM, nodes within the same community have the same expected degree while node degrees always vary in real-world networks because there are “hub” nodes with many more edges than other nodes in the same community. To solve this limitation, the Degree-Corrected Stochastic Blockmodel (DCSBM) DCSBM was proposed by introducing node heterogeneity parameters. Many works have been developed to study SBM and DCSBM, to name a few, rohe2011spectral ; lei2015consistency ; SCORE ; joseph2016impact ; abbe2016exact ; chen2018convexified . One limitation of SBM and DCSBM is, they assume that the network has non-overlapping properties while nodes can belong to multiple communities simultaneously in real-world social as well as other types of networks (for example, a movie can be classified into multiple genres, individuals belong to multiple social groups, scientists may belong to more than one research group, and proteins have multiple functions) palla2005uncovering ; xie2013overlapping ; javed2018community . To address this limitation, the Mixed Membership Stochastic Blockmodel (MMSB) MMSB and the Degree-Corrected Mixed Membership (DCMM) model MixedSCORE extend SBM and DCSBM to overlapping un-weighted networks, where DCMM can also be viewed as an extension of MMSB by considering node heterogeneity. In recent years, some spectral algorithms with theoretical guarantees on consistency have been proposed to study MMSB and DCMM, to name a few, MaoSVM ; mao2020estimating ; MixedSCORE . Thorough survey of the community detection approaches for un-weighted networks has been provided by jin2021survey .
@@
Edge weights are commonplace in complex networks and they provide more information for community detection power than edges newman2004analysis . Edge weights usually tells us how strong or weak a relationship is. For example, in a coauthorship network Newman2004 , edge weight between two authors is the number of papers they co-authored. In a signed network ferligoj1996analysis ; tang2016survey ; kirkley2019balance ; shi2019dynamics , positive edges represent friendship/trust and negative edges mean animosity/mistrust. Three examples of real-world weighted networks in which edge weights indicate the strength of relationship are shown in Figure 1. However, the aforementioned models ignore edge weights and they only model un-weighted networks. To address this limitation, some statistical models are developed in recent years. For non-overlapping weighted networks, several Weighted Stochastic Blockmodels (WSBMs) aicher2015learning ; jog2015information ; peixoto2018nonparametric ; palowitch2018significance ; xu2020optimal ; ng2021weighted , the Distribution-Free model (DFM) qing2021DFM , and the Degree-Corrected Distribution-Free model (DCDFM) DCDFM , and the bipartite versions of DFM and DCDFM BiDFMs extend SBM and DCSBM to weighted networks by allowing edge weights to follow distributions more than Bernoulli distribution. However, the above models for weighted networks assume that each node belongs to a single community and they can not capture the community structure of overlapping weighted networks. Though the weighted version of MMSB (WMMSB) dulac2020mixed and the multi-way blockmodels (MWB) airoldi2013multi can model overlapping weighted networks, they both face some limitations. WMMSB requires edge weights to be generated from Poisson distribution, so it fails to model un-weighted networks, signed networks, networks with positive and negative edge weights, and networks with decimal edge weights. MWB requires edge weights to follow Normal or Bernoulli distributions, thus it fails to describe the community structure of signed networks and networks in which edge weights are nonnegative integers. Meanwhile, both WMMSB and MWB ignore the variation of node degrees. To close these limitations of WMMSB and MWB, we aim at providing a new model for overlapping weighted networks. The main contributions of this article are summarized as follows.
(a) Slovene Parliamentary Party ferligoj1996analysis
(b) Gahuku-Gama subtribes read1954cultures
(c) Karate-club-weighted zachary1977information
@@
  • 1.
    We provide a generative model called Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model for overlapping networks with weighted edges. DCMMDF can be seen as an extension of some previous models. In particular, the aforementioned model DCMM is a sub-model of DCMMDF and overlapping signed networks can also be modeled by DCMMDF.
  • 2.
    We use a spectral algorithm with a theoretical guarantee of consistency to fit DCMMDF. The theoretical error rate for the algorithm when edge weights follow a specific distribution can be got immediately from the main result and we also provide several examples to show this.
  • 3.
    We provide overlapping weighted modularity in Equation (7) to measure the quality of overlapping community detection for weighted networks with positive and negative edge weights.
  • 4.
@@
    We implement experiments on computer-generated overlapping weighted networks and eleven real-world networks to demonstrate the advantages of the proposed model and the overlapping weighted modularity.
@@
The rest of the paper is organized as follows. Section 2 introduces the model. Section 3 introduces the algorithm fitting the model. Section 4 presents consistency of the algorithm under the model and provides some examples for different distributions. Section 5 introduces the overlapping weighted modularity and the method to determine the number of communities. Section 6 conducts extensive experiments. Section 7 concludes.
Notations. We use the following notations in this article. Write $[m]:=\{1,2,\ldots,m\}$ for any positive integer $m$. For a vector $x$ and fixed $q>0$, $\|x\|_{q}$ means its $l_{q}$-norm. For a matrix $M$, $M^{\prime}$ means its transpose, $\|M\|_{F}$ means its Frobenius norm, and $\|M\|_{2\rightarrow\infty}$ means the maximum $l_{2}$-norm among all rows of $M$. Let $\mathrm{rank}(M)$ be $M$’s rank. Let $\sigma_{k}(M)$ be $M$’s $k$-th largest singular value, $\lambda_{k}(M)$ be $M$’s $k$-th largest eigenvalue in magnitude, and $\kappa(M)$ be $M$’s condition number. Let $M(i,:)$ (and $M(:,j)$) be the $i$-th row (and the $j$-th column) of matrix $M$. $M(S_{r},:)$ and $M(:,S_{c})$ denote the rows and columns in the index sets $S_{r}$ and $S_{c}$ of matrix $M$, respectively. For any matrix $M$, let $Y=\mathrm{max}(0,M)$ be the matrix such that $Y_{ij}=\mathrm{max}(0,M_{ij})$ for any $i,j$. Let $e_{i}$ be the indicator vector with a $1$ in entry $i$ and 00 in all others.
2 The Degree-Corrected Mixed Membership Distribution-Free model
@@
Given a weighted network $\mathcal{N}$ with $n$ nodes (network is also known as graph and node is also known as a vertex in literature), let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix such that $A(i,j)$ denotes the edge weight between nodes $i$ and $j$ for $i,j\in[n]$. Since we only consider un-directed networks in this article, $A$ is symmetric. All entries of $A$ are allowed to be any finite real values in this article. We assume there are $K$ communities
@@
where node $i$ is a pure node if one element of $\Pi(i,:)$ is 1 and it is a mixed node otherwise. Let $\mathcal{I}$ be an index set of nodes corresponding to $K$ pure nodes, one from each community. Similar to mao2020estimating , we let $\Pi(\mathcal{I},:)$ be the $K\times K$ identity matrix for convenience. Let $P$ be a $K\times K$ block matrix (we also call $P$ connectivity matrix occasionally in this article) and satisfy the following conditions:
$P=P^{\prime},P\in\mathbb{R}^{K\times K},~{}\mathrm{rank}(P)=K,\mathrm{and~{}}P\mathrm{~{}has~{}unit~{}diagonals}.$ (4)
Note that since we consider weighted networks and $A$’s entries can by any finite real values, $P$ can have negative elements and $P$ is not a probability matrix like that of traditional models SBM SBM , DCSBM DCSBM , MMSB MMSB , and DCMM MixedSCORE for un-weighted networks. For convenience, let $P_{\mathrm{max}}=\mathrm{max}_{k,l\in[K]}|P(k,l)|$. Note that $P_{\mathrm{max}}$ can be larger than 1 because we will consider node heterogeneity later. Unless specified, throughout this article, $K$ is assumed to be a known integer.
@@
Equation (6) means that $A$’s expectation matrix $\Omega$ should have a block structure $\Theta\Pi P\Pi^{\prime}\Theta$ reflecting community membership information, and this is the reason we call $P$ a block matrix. Meanwhile, Equation (6) also means that $A$ is a random matrix generated from any distribution $\mathcal{F}$ with expectation matrix $\Omega$ and there is no limitation on $\mathcal{F}$ as long as Equation (6) holds.
@@
Call Equations (1)-(6) the Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model and denote it by $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$.
{tikzpicture}
[…]
{tikzpicture}
[…]
{tikzpicture}
[…]
{tikzpicture}
[…]
Figure 2: Illustrations for overlapping (and non-overlapping) networks can be modeled by DCMMDF. Panel (a): un-weighted network. Panel (b): weighted network with weights being nonnegative integers. Panel (c): weighted network with positive and negative weights. Panel (d): signed network.

### v3

# Overlapping community detection in weighted networks

@@ Abstract
Over the past decade, community detection in overlapping un-weighted networks, where nodes can belong to multiple communities, has been one of the most popular topics in modern network science. However, community detection in overlapping weighted networks, where edge weights can be any real value, remains challenging. In this article, we propose a generative model called the weighted degree-corrected mixed membership (WDCMM) model to model such weighted networks. This model adopts the same factorization for the expectation of the adjacency matrix as the previous degree-corrected mixed membership (DCMM) model. Our WDCMM extends the DCMM from un-weighted networks to weighted networks by allowing the elements of the adjacency matrix to be generated from distributions beyond Bernoulli. We first address the community membership estimation of the model by applying a spectral algorithm and establishing a theoretical guarantee of consistency. Then, we propose overlapping weighted modularity to measure the quality of overlapping community detection for both assortative and dis-assortative weighted networks. To determine the number of communities, we incorporate the algorithm into the proposed modularity. We demonstrate the advantages of the model and the modularity through applications to simulated data and real-world networks.
@@ 1 Introduction
Community detection is a popular data mining problem in network science and it is to partition a network into several components, where each component is called a community (also referred to as a module or a cluster, or a block). Generally, nodes within the same community have more interactions than nodes across communities. Community detection has many real-world applications in various domains, like sociology, biology, marketing, and computer science [1, 2, 3, 4, 5, 6, 7]. For example, community detection in biological networks helps researchers to analyze the interaction of brain regions and have a better understanding of the influence of the interaction on brain functions [7, 8, 9]. Community detection in social networks can help to identify crime organizations [10, 11]. Community detection in online social networks and food social networks can facilitate recommender systems [12, 13, 14]. Community detection in coauthorship and citation networks can be used to identify meaningful communities for scientists [15, 16].
@@
For its wide applications, in past decades, substantial methods have been developed to learn community structure for non-overlapping un-weighted networks in which each node belongs to a single community and the edges or links between nodes are either present or not [17, 18, 19], where un-weighted networks are also referred to as binary networks [20]. For community detection in non-overlapping un-weighted networks, perhaps the most popular and well-studied generative model is the Stochastic Blockmodel (SBM) [21]. The SBM assumes that the probability of generating an edge between two nodes depends on their belonging communities. Under SBM, nodes within the same community have the same expected degree while node degrees always vary in real-world networks because there are “hub” nodes with many more edges than other nodes in the same community. To solve this limitation, the Degree-Corrected Stochastic Blockmodel (DCSBM) [22] was proposed by introducing node heterogeneity parameters. Many works have been developed to study SBM and DCSBM, to name a few, [23, 24, 25, 26, 27, 28]. One limitation of SBM and DCSBM is, they assume that the network has non-overlapping properties while nodes can belong to multiple communities simultaneously in real-world social as well as other types of networks (for example, a movie can be classified into multiple genres, individuals belong to multiple social groups, scientists may belong to more than one research group, and proteins have multiple functions) [29, 30, 31]. To address this limitation, the Mixed Membership Stochastic Blockmodel (MMSB) [32] and the Degree-Corrected Mixed Membership (DCMM) model [33] extend SBM and DCSBM to overlapping un-weighted networks, where DCMM can also be viewed as an extension of MMSB by considering node heterogeneity and a reparametrization of the overlapping continuous community assignment model (OCCAM) considered in [34]. In recent years, some spectral algorithms with theoretical guarantees on consistency have been proposed to study MMSB and DCMM, to name a few, [35, 36, 37, 33]. A thorough survey of the community detection approaches for un-weighted networks has been provided by [38].
@@
Edge weights are commonplace in complex networks and they provide more information for community detection power than edges [39]. Edge weights usually tell us how strong or weak a relationship is. For example, in a coauthorship network [15], the edge weight between two authors is the number of papers they co-authored. In a signed network [40, 41, 42, 43], positive edges represent friendship/trust and negative edges mean animosity/mistrust. Three examples of real-world weighted networks in which edge weights indicate the strength of the relationship are shown in Figure 1. However, the aforementioned models ignore edge weights and they only model un-weighted networks. To address this limitation, some statistical models have been developed in recent years. For non-overlapping weighted networks, several Weighted Stochastic Blockmodels (WSBMs) [44, 45, 46, 47, 48, 49, 50, 51], the Distribution-Free model (DFM) [52], and the Degree-Corrected Distribution-Free model (DCDFM) [53], and the bipartite versions of DFM and DCDFM [54] extend SBM and DCSBM to weighted networks by allowing edge weights to follow distributions more than the Bernoulli distribution. Recently, [51] demonstrated that appropriate transformations of the adjacency matrix elements for a weighted network can be highly beneficial under WSBMs. Additionally, [55] also considered WSBMs that permit the adjacency matrix elements to follow any distribution and established the estimation consistency of a robust spectral clustering method for weighted networks. However, the above models for weighted networks assume that each node belongs to a single community and they can not capture the community structure of overlapping weighted networks. Though the weighted version of MMSB (WMMSB) [56] and the multi-way blockmodels (MWB) [57] can model overlapping weighted networks, they both face some limitations. WMMSB requires edge weights to be generated from Poisson distribution, so it fails to model un-weighted networks, signed networks, networks with positive and negative edge weights, and networks with decimal edge weights. MWB requires edge weights to follow Normal or Bernoulli distributions, thus it fails to describe the community structure of signed networks and networks in which edge weights are nonnegative integers. Meanwhile, both WMMSB and MWB ignore the variation of node degrees. To close these limitations of WMMSB and MWB, we aim to provide a model for overlapping weighted networks. The main contributions of this article are summarized as follows.
(a) Slovene Parliamentary Party [40]
(b) Gahuku-Gama subtribes [58]
(c) Karate-club-weighted [59]
@@
  • We provide a generative model called the Weighted Degree-Corrected Mixed Membership (WDCMM) model for overlapping networks with weighted edges. WDCMM shares the same factorization for the expectation of the adjacency matrix as DCMM and extends its applicability. This extension allows the elements of the adjacency matrix to be generated from various distributions, such as Normal, Poisson, and Uniform, in addition to the Bernoulli distribution. Moreover, signed networks can also be modeled by WDCMM. Further examples of distributions that WDCMM can handle are listed in Table 2.
  • We use a spectral algorithm with a theoretical guarantee of consistency to fit WDCMM. The theoretical error rate for the algorithm when edge weights follow a specific distribution can be obtained immediately from the main result and we also provide several examples to show this.
  • We provide overlapping weighted modularity in Equation (13) to measure the quality of overlapping community detection for both assortative and dis-assortative weighted networks with positive and negative edge weights.
@@
  • We implement experiments on computer-generated overlapping weighted networks and real-world networks to demonstrate the advantages of the proposed model and the overlapping weighted modularity.
@@
The rest of the paper is organized as follows. Section 2 introduces the model. Section 3 introduces the algorithm fitting the model. Section 4 presents the consistency of the algorithm under the model and provides some examples for different distributions. Section 5 introduces the overlapping weighted modularity and the method to determine the number of communities. Section 6 conducts extensive experiments. Section 7 concludes. All proofs are in the Appendix.
2 The Weighted Degree-Corrected Mixed Membership model
Symbol Description Symbol Description
$\mathbb{R}$ Real number set $\mathbb{R}_{+}$ Set of nonnegative real numbers
$\mathcal{N}$ Weighted network $P\in\mathbb{R}^{K\times K}$ Block matrix ($P=P^{\prime}$)
$K$ Number of communities $\|x\|_{q}$ $\ell_{q}$-norm for vector $x$
$n$ Number of nodes $^{\prime} XX$’s transpose
$[m]$ $\{1,2,\ldots,m\}$ for positive integer $m$ $\mathbb{E}$ Expectation
$A\in\mathbb{R}^{n\times n}$ Adjacency matrix $\|X\|_{2\rightarrow\infty}$ $X$’s maximum $\ell_{2}$-norm
$\mathcal{C}_{(k)}$ $k$-th community $\|X\|_{F}$ Frobenius norm of $X$
$\Pi\in[0,1]^{n\times K}$ Membership matrix $X(i,:)$ $X$’s $i$-th row
$\mathcal{I}$ Index set of pure nodes $X(:,j)$ $X$’s $j$-th column
$\rho$ Sparsity parameter $X(\mathcal{I},:)$ Rows in the index set $\mathcal{I}$ of $X$
$\mathcal{F}$ Distribution $\mathrm{rank}(X)$ $X$’s rank
$\theta$ $n\times 1$ node heterogeneity vector $\Theta$ Diagonal matrix with $\Theta(i,i)=\theta(i)$ for $i\in[n]$
$\theta_{\mathrm{max}}$ Maximum element of $\theta$ $\theta_{\mathrm{min}}$ Minimum element of $\theta$
$\sigma_{k}(X)$ $X$’s $k$-th largest singular value $\lambda_{k}(X)$ $X$’s $k$-th largest eigenvalue in magnitude
$\Omega$ $\Theta\Pi P\Pi^{\prime}\Theta$ $\kappa(X)$ $X$’s condition number
$P_{\mathrm{max}}$ Maximum entry of $P$ $\Omega_{\mathrm{max}}$ Maximum entry of $\Omega$
$U\in\mathbb{R}^{n\times K}$ Top $K$ eigenvectors of $\Omega$ $\mathcal{P}\in\{0,1\}^{K\times K}$ Permutation matrix
$\hat{U}\in\mathbb{R}^{n\times K}$ Top $K$ eigenvectors of AA $|a|$ Absolute value for real value $a$
$\Lambda$ Diagonal matrix with $\Lambda(k,k)=\lambda_{k}(\Omega)$ for $k\in[K]$ $\hat{\Lambda}$ Diagonal matrix with $\hat{\Lambda}(k,k)=\lambda_{k}(A)$ for $k\in[K]$
$_{*} UU$’s row-normalization version $\hat{U}_{*}$ $\hat{U}$’s row-normalization version
$\mathbb{N}$ Set of nonnegative integers $e_{i}$ $e_{i}(j)=1(i=j)$
$\hat{\Pi}\in[0,1]^{n\times K}$ Estimated memberships $\mathbb{P}$ Probability
$\tau$ $\mathrm{max}_{i\in[n],j\in[n]}|A(i,j)-\Omega(i,j)|$ $\gamma$ $\mathrm{max}_{i\in[n],j\in[n]}\frac{\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]}{\theta(i)\theta(j)}$
$\mathrm{diag}(M)$ Diagonal matrix with $(i,i)$-th entry $M(i,i)$ $\mathrm{max}(0,M)$ Matrix with $(i,j)$-th entry $\mathrm{max}(0,M(i,j))$
$M^{-1}$ Inverse of matrix $M$ $I_{K}$ $K\times K$ identity matrix
$Q_{ovw}$ Overlapping weighted modularity $n_{0}$ Number of pure nodes in each community
$\zeta_{\mathrm{mixed}}$ Proportion of highly mixed nodes $\zeta_{\mathrm{pure}}$ Proportion of highly pure nodes
$\varsigma$ Balance parameter $\hat{\mathcal{I}}$ Estimated index set
$\eta_{\mathcal{N}}$ Index reflecting $\mathcal{N}$’s assortativity $N_{M}$ Diagonal matrix with positive diagonal elements
$\mathbb{N}_{+}$ Set of positive integers $\mathbb{R}_{+}$ Set of positive real numbers
Table 1: Table of main symbols.
@@ 2 The Weighted Degree-Corrected Mixed Membership model
Table 1 summarizes the main symbols used in this paper. Given a weighted network $\mathcal{N}$ with $n$ nodes (network is also known as graph and node is also known as a vertex in literature), let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix such that $A(i,j)$ denotes the edge weight between nodes $i$ and $j$ for $i,j\in[n]$. Since we only consider un-directed networks in this article, $A$ is symmetric. All entries of $A$ are allowed to be any finite real values in this article. We assume there are $K$ communities
@@
where node $i$ is a pure node if one element of $\Pi(i,:)$ is 1 and it is a mixed node otherwise. Let $\mathcal{I}$ be an index set of nodes corresponding to $K$ pure nodes, one from each community, i.e., $\mathcal{I}=\{s_{1},s_{2},\ldots,s_{K}\}$, where $s_{k}$ is a pure node in the $k$-th community $\mathcal{C}_{(k)}$ for $k\in[K]$. Similar to [36], we let $\Pi(\mathcal{I},:)$ be the $K\times K$ identity matrix for convenience. Let $P$ be a $K\times K$ block matrix (we also call $P$ connectivity matrix occasionally in this article) and satisfy the following conditions:
$P=P^{\prime},P\in\mathbb{R}^{K\times K},~\mathrm{rank}(P)=K,\mathrm{and~}P\mathrm{~has~unit~diagonals},$ (4)
where the requirement for $P$ to have diagonal elements is imposed to ensure the identifiability of the proposed model. For further details, please refer to Remark 1 provided later. Note that since we consider weighted networks and $A$’s entries can by any finite real values, $P$ can have negative elements and $P$ is not a probability matrix like that of traditional models SBM [21], DCSBM [22], MMSB [32], and DCMM [33] for un-weighted networks. For convenience, let $P_{\mathrm{max}}=\mathrm{max}_{k,l\in[K]}|P(k,l)|$. Note that $P_{\mathrm{max}}$ can be larger than 1 because we will consider node heterogeneity later. Unless specified, throughout this article, $K$ is assumed to be a known integer.
@@
Equation (6) means that $A$’s expectation matrix $\Omega$ should have a block structure $\Theta\Pi P\Pi^{\prime}\Theta$ reflecting community membership information, and this is the reason we call $P$ a block matrix. Meanwhile, Equation (6) also means that $A$ is a random matrix generated from any distribution $\mathcal{F}$ with expectation matrix $\Omega$ and there is no limitation on $\mathcal{F}$ as long as Equation (6) holds. Furthermore, the expectation adjacency matrix $\Omega$ shares the same factorization as the $\Omega$ in Equation (2.3) of [33] under the DCMM model.
@@
Call Equations (1)-(6) the Weighted Degree-Corrected Mixed Membership (WDCMM) model and denote it by $WDCMM_{n}(K,P,\Pi,\Theta,\mathcal{F})$.
In Definition 1, we do not constrain any special type of the distribution $\mathcal{F}$ that can be handled by the WDCMM model. This is because our WDCMM allows $A$’s elements to be generated from any distribution $\mathcal{F}$ as long as $A$’s expectation under $\mathcal{F}$ is $\Omega$, which enjoys a block structure related to nodes’ mixed memberships. In other words, the only limitation of our WDCMM model on the distribution $\mathcal{F}$ is that $\mathbb{E}[A(i,j)]=\Omega(i,j)$ for $i,j\in[n]$ under the distribution $\mathcal{F}$, i.e., $\mathcal{F}$’s first moment exists and has a block structure $\Omega$. For example, $\mathcal{F}$ can be distributions like Normal, Bernoulli, Poisson, Uniform, Binomial, Negative binomial, Exponential, Gamma, Beta, Geometric, Laplace, Lognormal, Pareto, Logistic, and Weibull distributions as described in [60], as well as a discrete distribution for signed networks. The details of these distributions are provided below, for $i,j\in[n]$:
  • $\mathcal{F}$ can be Normal distribution by letting $A(i,j)\sim\text{Normal}(\Omega(i,j),\sigma^{2}_{A})$. For this distribution, $A(i,j)\in\mathbb{R}$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and there is no limitation on $\Omega(i,j)$.
  • $\mathcal{F}$ can be Bernoulli distribution by letting $A(i,j)\sim\text{Bernoulli}(\Omega(i,j))$. For this distribution, $A(i,j)\in\{0,1\}$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be a probability in $[0,1]$.
  • $\mathcal{F}$ can be Poisson distribution by letting $A(i,j)\sim\text{Poisson}(\Omega(i,j))$. For this distribution, $A(i,j)$ is a nonnegative integer, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be a nonnegative value.
  • $\mathcal{F}$ can be Uniform distribution by letting $A(i,j)\sim\text{Uniform}(a,2\Omega(i,j)-a)$ (or $A(i,j)\sim\text{Uniform}(2\Omega(i,j)-a,a)$) for any finite value $a$. For this distribution, $A(i,j)\in(a,2\Omega(i,j)-a)$ (or $A(i,j)\in(2\Omega(i,j)-a,a)$), $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be larger (or smaller) than $a$.
  • $\mathcal{F}$ can also be a discrete distribution such that $P(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $P(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$. For this case, $A$ is the adjacency matrix of a signed network, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should range in [1,1][-1,1].
  • $\mathcal{F}$ can be Binomial distribution by letting $A(i,j)\sim\text{Binomial}(m,\frac{\Omega(i,j)}{m})$, where $m$ is a positive integer. For this distribution, $A(i,j)$ is a nonnegative integer that is no larger than $m$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be a value ranging in $[0,m]$ since $\frac{\Omega(i,j)}{m}$ is a probability.
  • $\mathcal{F}$ can be Negative binomial distribution by letting $A(i,j)\sim\text{Negative binomial}(m,\frac{m}{m+\Omega(i,j)})$, where $m$ is a positive integer. For this distribution, $A(i,j)$ is a nonnegative integer, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be nonnegative.
  • $\mathcal{F}$ can be Exponential distribution by letting $A(i,j)\sim\text{Exponential}(\frac{1}{\Omega(i,j)})$. For this distribution, $A(i,j)$ is nonnegative, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be positive.
  • $\mathcal{F}$ can be Gamma distribution by letting $A(i,j)\sim\text{Gamma}(a,\frac{\Omega(i,j)}{a})$ (or $A(i,j)\sim\text{Gamma}(\frac{\Omega(i,j)}{a},a)$) for any $a>0$. For this distribution, $A(i,j)$ is positive, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be positive.
  • $\mathcal{F}$ can be Beta distribution by letting $A(i,j)\sim\text{Beta}(a\frac{\Omega(i,j)}{1-\Omega(i,j)},a)$ (or $A(i,j)\sim\text{Beta}(a,a\frac{1-\Omega(i,j)}{\Omega(i,j)})$) for any $a>0$. For this distribution, $A(i,j)\in(0,1)$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should range in $(0,1)$.
  • $\mathcal{F}$ can be Geometric distribution by letting $A(i,j)\sim\text{Geometric}(\frac{1}{\Omega(i,j)})$. For this distribution, $A(i,j)$ is a positive integer, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be no smaller than 1 given that $\frac{1}{\Omega(i,j)}$ is a probability.
  • $\mathcal{F}$ can be Laplace distribution (also known as Double exponential distribution) by letting $A(i,j)\sim\text{Laplace}(\Omega(i,j),a)$ for any $a>0$. For this distribution, $A(i,j)\in\mathbb{R}$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ can be any finite value.
  • $\mathcal{F}$ can be Lognormal distribution by letting $A(i,j)\sim\text{Lognormal }(\mathrm{log}(\Omega(i,j))-\frac{a^{2}}{2},a^{2})$ for $a>0$. For this distribution, $A(i,j)>0$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be positive.
  • $\mathcal{F}$ can be Pareto distribution by letting $A(i,j)\sim\text{Pareto}(\frac{a-1}{a}\Omega(i,j),a)$ for $a>2$. For this distribution, $A(i,j)>\frac{a-1}{a}\Omega(i,j)$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be positive.
  • $\mathcal{F}$ can be Logistic distribution by letting $A(i,j)\sim\text{Logistic}(\Omega(i,j),a)$ for any $a>0$. For this distribution, $A(i,j)\in\mathbb{R}$, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ can be any finite value.
  • $\mathcal{F}$ can be Weibull distribution by letting $A(i,j)\sim\text{Weibull}(a,(\frac{\Omega(i,j)}{\Gamma(1+\frac{1}{a})})^{a})$ for $a>0$, where $\Gamma(\cdot)$ is the gamma function. For this distribution, $A(i,j)$ is nonnegative, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds, and $\Omega(i,j)$ should be positive.
Certainly, not all distributions satisfy Equation (6). Below are some examples:
  • $\mathcal{F}$ cannot be a degenerate distribution with parameter $a$, where the random variable is constant and equals $a$ with probability 1.
  • $\mathcal{F}$ cannot be a t-distribution whose mean is 0.
  • $\mathcal{F}$ cannot be a Cauchy distribution, as its mean does not exist.
  • $\mathcal{F}$ cannot be a discrete uniform distribution on the first $m$ natural numbers because its expectation, $\frac{m+1}{2}$, is a fixed value that cannot capture the block structure in $\Omega$.
  • Given that the Hypergeometric distribution describes the probability of obtaining exactly $m$ successes in a sample of $N_{2}$ items drawn without replacement from a population of $N$ items, where $N_{1}$ items are considered successes, its mean, $\frac{N_{1}N_{2}}{N}$, cannot capture the block structure in $\Omega$. Therefore, $\mathcal{F}$ cannot be a Hypergeometric distribution. Similarly, $\mathcal{F}$ cannot be a Negative-Hypergeometric distribution.
  • Given that the mean of a Chi-square distribution ($\chi^{2}$-distribution) with $m$ degrees of freedom is $m$, a positive integer than cannot capture the block structure in $\Omega$. Thus, $\mathcal{F}$ cannot be a $\chi^{2}$-distribution.